OpenStudy (mallorysipp234):

???? 3x^-2+2x^-1-8=0 How do i find all solutions to this equation?

OpenStudy (mallorysipp234):

i've never worked with negative exponents in an equation like this and dont know how to begin working with it.

OpenStudy (s4sensitiveandshy):

There are two ways to solve this question: first) move the constant term to the right side and then take out the common factor 2nd) change negative exponent to positive exponents and then find the common denominator

OpenStudy (mallorysipp234):

Isn't the constant term and common factor the same thing in this scenario?

OpenStudy (agent0smith):

Multiply both sides by x^2, that'll positivize all the exponents.

OpenStudy (s4sensitiveandshy):

Hmm i don't think so. the constant term isn't the common factor.

OpenStudy (s4sensitiveandshy):

*multiply by x^2* that is a good approach suggested by agent one step and then you will get the quadratic equation. go with that.

OpenStudy (anonymous):

put \[x ^{-1}=y,squaring ~,\left( x ^{-1} \right)^2=y^2~or~x ^{-2}=y^2\] \[3y^2+2y-8=0\] \[3y^2+6y-4y-8=0\] make factors and find y then \[y=x ^{-1}=\frac{ 1 }{ x },~or~x=\frac{ 1 }{ y }\]

OpenStudy (anonymous):

\[\frac{3}{x^2}+\frac{2}{x}-8=0 \]\[\left\{x=-\frac{1}{2},x=\frac{3}{4}\right\} \]

OpenStudy (mallorysipp234):

@robtobey so initially dividing the terms by their power and then solving for x works? all the same?

OpenStudy (agent0smith):

He didn't divide the terms by their power... \[\Large x^{-a} = \frac{ 1 }{ x^a }\]